How to substute value like 3a+4a when a=5 and 5g*g when g=30 

Grade:7

2 Answers

Arun
4 years ago
Dear student
 
3a + 4a = 7a
Now put a = 5
Hence 7a = 7*5 = 35
 
Similarly
 
5g *g = 5*30*30 = 4500
 
Hope it helps
 
In case of any difficulty, please feel free to ask
Yashraj
3 years ago
please clear your question attach an image and try again later...................................................................................